OpenStudy (anonymous):

Simplify the problem below. -5 × 9 -------------- -21 + 6 Select one of the options below as your answer: A. -3 B. - 5 3 C. 5 3 D. 3

OpenStudy (kaiz122):

just do simple multiplication, addition and division. evaluate the numerator and denominator and you will come up with the answer

OpenStudy (kaiz122):

\[\frac{-5*9}{-21+6} = \frac{-45}{-15}\]

OpenStudy (anonymous):

Okay This Is What I Did